# Sample size and power calculations for causal inference in observational studies

### v3

@@ 4.1 Synthetic Data
We also conduct simulation studies where (i) the outcomes are binary and (ii) the estimand is the average treatment effect for the treated or the overlap population, the details of which are relegated to Supplementary Material D.3 and D.4, respectively. The proposed method is shown to correctly estimate the sample size under both settings.
4.2 Real data based simulations
@@
This section illustrates the proposed methods using simulations based on a widely used benchmark data from the Right Heart Catheterization observational study (Connors et al.,, 1996), which was conducted to evaluate the causal effect of the diagnostic cardiological procedure of right heart catheterization (binary treatment) among hospitalized critically ill patients. The purpose of our simulation is not to re-design an existing study; instead, we want to assess the performance of our methods in a real world setting where we do not know the true data generating model, presenting a more realistic and challenging setting. The data contain 5735 patients, each with 51 covariates (20 continuous, 25 binary, 6 categorical). The outcome is the binary survival status at 30 days after admission, with $Y=1$ meaning death. In total, 2184 ($38.1\%$) of the patients received the treatment and the rest did not. The death rate is 38.0% and 30.6% in the treatment and control group, respectively.
@@
The simulation proceeds as follows. We first calculate the summary parameters $(r,\phi,\rho^{2})$ and effect size $\widetilde{\tau}^{\mbox{\tiny{ATE}}}$ from the real data. We take these values as the only inputs of a hypothetical power analysis, without access to the whole data. We apply the proposed method with these inputs to compute the minimal sample size needed to achieve a certain level of power in detecting the pre-fixed effect size $\widetilde{\tau}^{\mbox{\tiny{ATE}}}$. We then reverse the process to estimate the power under a sample of that calculated size. The closer the estimated power is to the nominal power (e.g. 0.8), the more accurate the our method is. We did not compare with the method of Shook-Sa and Hudgens, (2022) because it requires the distribution of weights, which is assumed not available in this study.
@@
Following the above outline, from the real data we estimate the propensity scores using a logistic model with the main effects of each covariate (the distributions by group are shown in Supplementary Material C.1 (Figure 6). Based on these estimated scores, we numerically compute the true overlap coefficient $\rho$ defined in (9) to be 0.835. We also used the proposed normal approximation in Section 3.2, $W_{e}\sim\mathsf{N}(-0.701,2.189)$, and estimate the overlap coefficient to be 0.822, which is close to the true value, showing the precision of our proposed Beta approximation. The Hájek estimate of the standardized effect size $\widetilde{\tau}^{\mbox{\tiny{ATE}}}$ is 0.14. The treatment proportion is $r=0.381$ and $\rho^{2}=0.000$. With these inputs alone, we apply formula (17) to calculate the minimal sample size to achieve 0.8 power to be 3625. We then reverse the process to calculate the actual power under a sample of 3625 units. Specifically, we generate $B=1000$ replicates, each with 3625 rows randomly drawn from the original data with replacement. We calculate the Hájek estimate of $\tau^{\mbox{\tiny{ATE}}}$ and the associated $95\%$ confidence interval in each replicate. Among the 1000 replicates, 924 intervals do not cover zero, giving an estimated power of 0.924 with 95% confidence interval (0.908 – 0.940), which is larger than the nominal power of 0.8. Potential reasons for the discrepancy include: (i) we used the estimated instead of the true propensity scores, and (ii) we treated the binary outcome as Gaussian. We also computed the sample size calculated by the two-sample $z$-test, which is 1567, leading to a vastly underpowered estimate with the power of only 0.55. Given that any real study would have data attrition due to missing data and extreme weights, a conservative (overpowered) estimate of the sample size is arguably more desirable than the opposite.
@@
Figure 3 shows the power curves of $\tau^{\mbox{\tiny{ATE}}}$ given different values of the overlap coefficient $\phi$ and the correlation $\rho_{1}^{2}=\rho_{0}^{2}=\rho^{2}$ while fixing other summary inputs as those calculated from the Right Heart Catheterization data. It shows that, to achieve the same power, the minimal sample size would increase as either the overlap decreases or the correlation $\rho$ increases. Both patterns are as expected from the theory. The $R^{2}$ value is 0.21, which serves as the upper bound of $\rho^{2}$.
We repeat the process for the average treatment effect for the treated and the overlap population, respectively. The results are relegated to Supplementary Material D.5.
@@ 5 Discussion
This paper investigates the theoretical foundation for power analysis in the propensity score analysis of causal inference with observational data. We show that, compared with the power calculation of randomized trials, users only need to specify two parameters for emulated trials, which quantifies the strength of the confounder-treatment and the confounder-outcome association, respectively. We propose interpretable and easily computable statistics to characterize these two parameters. Our method bypasses distributional assumptions on the multivariate covariates and generalizes the existing variance inflation factor method. We provide new theoretical insights on the sufficient summary information for power calculation in observational studies.
@@
There are several important extensions of our method in observational studies, including repeated outcomes (Liu and Wu,, 2005), time-to-event outcomes (Hsieh and Lavori,, 2000), and multilevel data (Arpino and Mealli,, 2011). Propensity weighting estimators for causal effects are available in all these settings. Therefore, we can readily adapt the general structure of power calculation developed in this paper. For example, for studies with repeated outcomes, a common estimand is the time-averaged difference; we can analyze the variance of an inverse probability weighted marginal outcome model within-subject correlation between outcomes. For time-to-event outcomes, a common estimand is the marginal causal hazard ratio from the Cox proportional hazard model; we can analyze the variance of the inverse probability weighted Cox model.
Supplementary Materials
A Proof of Propositions and Theorems
A.1 Proof of (5)
@@
$\Pr\left(\tau+\epsilon\sqrt{\frac{V_{0}}{N}}>z_{1-\alpha}\sqrt{\frac{V}{N}}\right)$ $=\Pr\left(\epsilon>z_{1-\alpha}\sqrt{\frac{V}{V_{0}}}-\tau\sqrt{\frac{N}{V_{0}}}\right)$ (A.1)
$=1-\Phi\left(\frac{z_{1-\alpha}\sqrt{V}-\tau\sqrt{N}}{\sqrt{V_{0}}}\right).$ (A.2)
@@
$N=\frac{\left(z_{1-\alpha}\sqrt{V}-z_{1-\beta}\sqrt{V_{0}}\right)^{2}}{\tau^{2}}.$ (A.3)
@@
$N\leq\frac{V}{\tau^{2}}\left(z_{1-\alpha}-z_{1-\beta}\right)^{2}=\frac{V}{\tau^{2}}\left(z_{1-\alpha}+z_{\beta}\right)^{2}.$ (A.4)
@@ A.2 Special case of the Lyapunov Central Limit Theorem
$\lim_{n\rightarrow\infty}\frac{1}{s_{n}^{2+\delta}}\sum_{i=1}^{n}\mathbb{E}\left[|Z_{i}-\mu_{i}|^{2+\delta}\right]=0$ (A.5)
@@
$\frac{1}{s_{n}}\sum_{i=1}^{n}(Z_{i}-\mu_{i})\Rightarrow\mathsf{N}(0,1).$ (A.6)
@@ A.3 Proof of Proposition 1
$\mathbb{E}[Z]=\mathbb{E}[\mathbb{E}[Z\mid U]]=\mathbb{E}[U]=\frac{a}{a+b}.$ (A.7)

### v4

@@ 5.1 Synthetic Data: homogeneous treatment effect
We also conduct simulation studies where (i) the outcomes are binary and (ii) the estimand is the ATT or ATO, details of which are relegated to Appendix C.1 and C.2, respectively. The proposed method is shown to correctly estimate the sample size in both settings.
5.2 Synthetic Data: heterogeneous treatment effect
We extend the simulation in Section 5.1 to assess the proposed method in the presence of heterogeneous treatment effect (HTE). We generate covariates, propensity score, treatment indicator, and the control potential outcome $Y_{i}(0)$ following Section 5.1, and simulate $N=1,000,000$ units as the superpopulation. Following the treatment effect decomposition in Ding et al. (2019), we generate the individual treatment effect as $\tau_{i}=\tau_{0}+g(X_{i})+u_{i}$, where $g(X_{i})$ is the systematic component driven by observed covariates and $u_{i}\sim\mathsf{N}(0,0.2^{2})$ is the idiosyncratic component independent of observed covariates. The treated potential outcome is $Y_{i}(1)=Y_{i}(0)+\tau_{i}$. We target the ATE estimand $\tau=\operatorname{\mathbb{E}}[\tau_{i}]$. We consider two systematic components: (i) covariate-driven HTE, where $g(X_{i})=\delta(X_{1,i}+X_{9,i})$, so that treatment effect heterogeneity is driven by both a confounder and a precision variable, and (ii) propensity score-driven HTE, where $g(X_{i})=\delta W_{e,i}$, so that treatment effect varies along the logit propensity score. For each scenario, we calibrate $(\tau_{0},\delta)$ so that $\operatorname{\mathbb{E}}[\tau_{i}]=\operatorname{\mathbb{V}}(\tau_{i})=1$ in the superpopulation, representing substantial effect heterogeneity where the standard deviation of individual treatment effects equals the true ATE. $\phi=1$ ($\kappa=0$) is not applicable to propensity score-driven HTE because $W_{e}$ is constant, and is thus omitted. By construction, this HTE mechanism increases $\operatorname{\mathbb{V}}[Y(1)]$ while $\operatorname{\mathbb{V}}[Y(0)]$ remains unchanged, a feature also shared by control potential outcome-driven HTE simulation designs in Ding et al. (2015); Branson and Dasgupta (2020).
As in Section 5.1, from the superpopulation, we estimate $W_{e}$ by $\widehat{W}_{e}$ with a logistic model and $\tau$ by $\hat{\tau}^{\mbox{\tiny{OW}}}$ with the overlap weight estimator. Under the homogeneity assumption, we impute all $Y_{i}(0)$ for the group $Z_{i}=1$ by $Y_{i}-\hat{\tau}^{\mbox{\tiny{OW}}}$ and all $Y_{i}(1)$ for the group $Z_{i}=0$ by $Y_{i}+\hat{\tau}^{\mbox{\tiny{OW}}}$. Then, we compute the pooled variance $S^{2}=\operatorname{\mathbb{V}}[Y(0)]$, correlation $\rho=\mathrm{cor}(Y(0),W_{e})$, and R-squared statistic $R^{2}=\operatorname{\mathbb{V}}[f(X)]/\operatorname{\mathbb{V}}[Y(0)]$ empirically. Although $\tau_{i}$ varies individually, this homogeneous imputation ensures empirical $S^{2}$, $\rho$, and $R^{2}$ calculated with imputed $Y(z)$ are the same for $z=0,1$ in the superpopulation. The $(r,\phi)$ remains unchanged from Section 5.1. Similarly, we use $(r,\phi,\rho^{2})$ and the standardized effect $\tilde{\tau}=1/S$ to calculate the required sample size $N_{\mathrm{pos}}$ and estimate the empirical power through $B=10,000$ replications. The empirical power is the proportion of rejecting the null hypothesis in test (4) at level $\alpha=0.05$ with the Hájek estimator over $B$ replications. We report the simulation results in Table 2. We also conduct simulations using the same simulated superpopulations under ATT and ATO estimands, and results are relegated to Appendix C.3.
Table 2: Sample size calculation by the proposed method, the two-sample $z$-test and the method in (Shook-Sa and Hudgens, 2022) to achieve nominal power of 0.80 under various degrees of overlap and HTE. The power is obtained from random sampling of the calculated size from the N=1,000, [^0] N=1{,}000{,}000 units. The Monte Carlo error for a power of 0.8 is 0.004. The empirical $S^{2}$ is about 21.1 and 19.5 for the upper and lower panels, respectively.
Proposed $z$-test Shook-Sa & Hudgens
$\phi$ $\rho^{2}$ $R^{2}$ Size True PS Est PS Size True PS Est PS Size True PS Est PS
covariates-driven HTE
1.00 0.02 0.23 663 0.82 0.89 663 0.82 0.89 663 0.81 0.88
0.98 0.03 0.23 693 0.80 0.88 664 0.80 0.88 691 0.82 0.88
0.93 0.03 0.24 795 0.81 0.88 665 0.72 0.81 787 0.81 0.88
0.87 0.03 0.24 1048 0.81 0.87 666 0.61 0.67 $10$ 0.78 0.84
0.83 0.03 0.24 1363 0.80 0.85 666 0.49 0.53 1253 0.75 0.81
0.81 0.03 0.24 1715 0.79 0.84 667 0.37 0.42 1528 0.75 0.74
propensity score-driven HTE
0.98 0.01 0.17 641 0.79 0.86 616 0.79 0.85 640 0.80 0.86
0.93 0.00 0.17 727 0.80 0.85 613 0.72 0.78 726 0.79 0.85
0.87 0.00 0.17 942 0.78 0.82 612 0.60 0.63 928 0.77 0.83
0.83 0.00 0.17 1207 0.78 0.82 611 0.49 0.47 1150 0.77 0.80
0.81 0.00 0.17 1499 0.73 0.78 611 0.39 0.37 1401 0.71 0.76
Table 2 reports the calculated sample sizes and empirical power under HTE. Remark 1 notes that under HTE, the asymptotic variance of the Hájek estimator additionally depends on $\operatorname{\mathbb{V}}[\tau(X)]$, which is generally unavailable at the design stage. Nevertheless, the proposed method continues to achieve empirical power close to the nominal 0.80 across almost all overlap levels and both HTE scenarios, even when $\operatorname{\mathbb{V}}(\tau_{i})$ equals the true ATE. Similar to the case of homogeneous effect, the power with estimated propensity scores exceeds, often by a sizeable margin, that with the true scores under HTE. The comparison with the $z$-test and the method of Shook-Sa and Hudgens (2022) mirrors the homogeneous case: the $z$-test increasingly underestimates the required sample size as overlap deteriorates, while the SH method tracks the proposed method closely when $\rho^{2}$ is small, consistent with Remark 3.
5.3 Real-data-based nonparametric simulations
@@
This section illustrates the proposed methods using simulations based on a widely used benchmark data from the Right Heart Catheterization (RHC) observational study (Connors et al., 1996), which was conducted to evaluate the causal effect of the diagnostic cardiological procedure of RHC (binary treatment) among hospitalized critically ill patients. The purpose of our simulation is not to re-design an existing study or conduct a post-hoc power analysis; instead, we want to assess the performance of the proposed methods in a real world setting where we do not know the true data generating model, presenting a more realistic and challenging setting. The data contain 5735 patients, each with 51 covariates (20 continuous, 25 binary, 6 categorical). The outcome is the binary survival status at 30 days after admission, with $Y=1$ meaning death. In total, 2184 ($38.1\%$) of the patients received the treatment and the rest did not. The death rate is 38.0% and 30.6% in the treatment and control group, respectively.
@@
The simulation is outlined as follows. First, we calculate the summary parameters $(r,\phi,\rho^{2})$ and the effect size $\widetilde{\tau}^{\mbox{\tiny{ATE}}}$ from the real data. Second, we take these values as the only inputs of a hypothetical power analysis for a new observational study with the same goal on a similar population. We apply the proposed formula in Theorem 2 with these inputs to compute the minimal sample size $N$ needed to achieve a certain level of power in detecting the pre-fixed effect size $\widetilde{\tau}^{\mbox{\tiny{ATE}}}$. Third, we reverse the process to estimate the power using repeated bootstrapping samples of size $N$ randomly drawn from the RHC data. The closer the estimated power is to the nominal power (e.g. 0.8), the more accurate our method is. This simulation is nonparametric because we do not assume any outcome model.
@@
Following the above outline, from the real data we estimate the propensity scores using a logistic model with the main effects of each covariate. Based on these estimated scores, we numerically compute the true overlap coefficient $\phi$ defined in (9) to be 0.835. We also used the proposed normal approximation in Section 3.2, $W_{e}\sim\mathsf{N}(-0.701,2.189)$, and estimate the overlap coefficient to be 0.822, which is close to the true value, showing the precision of our proposed Beta approximation. The Hájek estimate of the standardized effect size $\widetilde{\tau}^{\mbox{\tiny{ATE}}}$ is 0.14. The treatment proportion is $r=0.381$ and $\rho^{2}=0.000$. With these inputs alone, we apply formula (14) to calculate the minimal sample size to achieve 0.8 power to be 3625. We then reverse the process to calculate the actual power under a sample of 3625 units. Specifically, we generate $B=1000$ replicates, each with 3625 rows randomly drawn from the original data with replacement. We calculate the Hájek estimate of $\tau^{\mbox{\tiny{ATE}}}$ and the associated $95\%$ confidence interval in each replicate. Among the 1000 replicates, 924 intervals do not cover zero, giving an estimated power of 0.924 with 95% confidence interval (0.908 – 0.940), which is larger than the nominal power of 0.8. Potential reasons for the discrepancy include: (i) we used the estimated instead of the true propensity scores, and (ii) we treated the binary outcome as Gaussian. We also computed the sample size calculated by the two-sample $z$-test, which is 1567, leading to a vastly underpowered estimate with the power of only 0.55. Given that any real study would have data attrition due to missing data and extreme weights, a conservative (overpowered) estimate of the sample size is arguably more desirable than the opposite.
@@
Figure 4 shows the power curves of $\tau^{\mbox{\tiny{ATE}}}$ given different values of the overlap coefficient $\phi$ and the correlation $\rho^{2}$ while fixing other summary inputs as those calculated from the RHC data. It shows that, to achieve the same power, the minimal sample size would increase as either the overlap $\phi$ decreases or the correlation $\rho$ increases, matching the theoretical result in Corollary 1. The $R^{2}$ value is 0.21, which serves as the upper bound of $\rho^{2}$.
We repeat the process for the ATO and ATT estimand, respectively, and the results suggest the proposed formula lead to reliable power calculation in a wide range of scenarios. Details are relegated to Appendix C.4.
@@ 6 Discussion
This paper establishes the theoretical foundation of power analysis for causal inference with observational data. We show that, compared with the power calculation of randomized trials, users only need to specify two additional parameters for emulated trials: an overlap coefficient $\phi$ that quantify the confounder-treatment association and a correlation parameter $\rho$ that quantify the confounder-outcome association. We propose interpretable and easily computable statistics to characterize these two parameters. Our method requires a logistic model for the propensity scores and a homoscedastic restricted mean outcome model with a homogeneous treatment effect, but it bypasses distributional assumptions on the multivariate covariates due to the Lyapunov-type central limit theorem. In practice, investigators can set these parameters based on a pilot study or similar previous studies, or conduct related sensitivity analysis based on domain knowledge of these parameters.
A power analysis must be conducted before data collection. Once a study is carried out, it is generally not meaningful and in fact is misleading to perform a post-hoc power analysis on the collected data, e.g. using the bootstrapping procedure in the RHC simulation in Section 5.3. Such a type of simulation cannot capture all the uncertainties in the causal effect because of the conditioning on the observed sample. Moreover, the observed data may not satisfy the model assumptions (e.g. effect homogeneity) underlying the sample size formula that the post-hoc power analysis implicitly assumes.
Some authors (e.g. Hernán, 2022) questioned the role of power analysis in observational studies; they argued that the primary consideration in observational studies is to reduce confounding bias rather than to achieve sufficient power. We agree on the central role of de-confounding in causal inference, but that does not preclude the critical role of power analysis in designing observational studies, which are usually subject to cost and time constraints. For example, in prospective studies of rare or severe diseases, recruiting patients is time-consuming, and thus investigators must determine the minimal sample size to achieve a reliable estimate before conducting the study. In this type of cohort studies, ideally investigators could recruit participants in batches, with the first batch serving as a pilot study, which provides reliable estimate of the key inputs $(\phi,\rho)$ and enables a more accurate calculation of the sample size needed for the subsequent batches. Another example concerns retrospective commercial population health data (such as the Truveta electronic health record data or IBM’s MarketScan health insurance claims data), which usually charge users costs proportional to the data size and computational load. With limited resources, a sample size calculation is crucial for designing studies based on such data sources. Moreover, the same data source (e.g. disease registries or electronic health records of a population) are often used for investigating multiple hypotheses, e.g. effects on secondary outcomes in clinical research. In these cases, power analysis can help researchers decide how aggressive they can be with multiplicity correction or if they have to whittle down the primary objective.
@@
There are several extensions of the proposed method in observational studies, including repeated outcomes, time-to-event outcomes, and multiple treatments. Propensity score weighting estimators for causal effects are available in all these settings. Therefore, we can readily adapt the general structure of power calculation developed in this paper. For example, for studies with repeated outcomes, a common estimand is the time-averaged difference; we can analyse the variance of an inverse probability weighted marginal outcome model within-subject correlation between outcomes. For time-to-event outcomes, a common estimand is the marginal causal hazard ratio from the Cox proportional hazard model; we can analyse the variance of the inverse probability weighted Cox model. Then, we can use the Beta approximation of the propensity score distribution to derive a closed-form expression of the variance inflation factor of the weighting estimator between an emulated trial and the corresponding randomized trial (Kish, 1965).
Acknowledgements
The authors thank the associate editor, three anonymous reviewers, Xiaoxiao Zhou, Jay Lusk, and Laine Thomas for constructive comments. This paper is dedicated to the memory of the late Sayan Mukherjee, who taught FL basics of power analysis and beyond.
Appendix
Appendix A Proof of Propositions and Theorems
A.1 Proof of Equation (5)
@@
$\Pr\left(\tau+\epsilon\sqrt{\frac{V_{0}}{N}}>z_{1-\alpha}\sqrt{\frac{V}{N}}\right)$ $=\Pr\left(\epsilon>z_{1-\alpha}\sqrt{\frac{V}{V_{0}}}-\tau\sqrt{\frac{N}{V_{0}}}\right)$ (21)
$=1-\Phi\left(\frac{z_{1-\alpha}\sqrt{V}-\tau\sqrt{N}}{\sqrt{V_{0}}}\right).$ (22)
@@
$N=\frac{\left(z_{1-\alpha}\sqrt{V}-z_{1-\beta}\sqrt{V_{0}}\right)^{2}}{\tau^{2}}.$ (23)
@@
$N\leq\frac{V}{\tau^{2}}\left(z_{1-\alpha}-z_{1-\beta}\right)^{2}=\frac{V}{\tau^{2}}\left(z_{1-\alpha}+z_{\beta}\right)^{2}.$ (24)
@@ A.2 Special case of the Lyapunov Central Limit Theorem
$\lim_{n\rightarrow\infty}\frac{1}{s_{n}^{2+\delta}}\sum_{i=1}^{n}\mathbb{E}\left[|Z_{i}-\mu_{i}|^{2+\delta}\right]=0$ (25)
@@
$\frac{1}{s_{n}}\sum_{i=1}^{n}(Z_{i}-\mu_{i})\Rightarrow\mathsf{N}(0,1).$ (26)
@@ A.3 Proof of Proposition 1(i)
$\mathbb{E}[Z]=\mathbb{E}[\mathbb{E}[Z\mid U]]=\mathbb{E}[U]=\frac{a}{a+b}.$ (27)